OpenStudy (anonymous):

Find the values of x,y and z (4z+2) (3y-11) (y+19)

OpenStudy (agreene):

there is no x. y= -19 y=11/3 z=-1/2 are the solutions.

OpenStudy (anonymous):

Thanks!